# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ 4.4 Hallucination Cost
Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $H\colon\mathcal{R}\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ measures discrepancy:
@@ 4.5.1 Truthfulness
The fundamental requirement for a reliable inference system is that its components do not misrepresent their knowledge. The principle of truthfulness, also known as incentive compatibility in mechanism design, ensures that each agent’s optimal strategy is to accurately represent its privately held informations.
@@
A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that are perfectly aligned with the accessible private knowledge. Any other alternative response $\bar{s}_{i}$ that is not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
@@
$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$
@@
With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately.
@@ 4.5.2 Semantic Information Conservation
Second, we want to consider mechanisms that cannot create knowledge ex nihilo. The Semantic Information Conservation principle demands that the net informational contribution across all agents sums to zero. This property is the formal guardrail against ungrounded fabrication and ensures that any generated response, no matter how creative, is ultimately derived from the system’s existing knowledge
@@
This principle constrains creative information generation, imagination if you will. The total information contributions must balance to zero, preventing fabrication of knowledge not derivable from available sources. That also leads to response with bounded creativity.
Definition 14 (Bounded Response Creativity).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Let $B^{*}_{C}$ be the fixed point of reasoning for the baseline knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (22)
The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$. Let us see why.
Theorem 5 (Bounded Creativity of Chain of Thoughts).
Let $q\in\mathcal{Q}$ be a query with context $Q=K(q)\cup T(q)$ and let $B^{*}_{C}$ be the fixed point of reasoning for the baseline context-relevant knowledge $B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$. Let us assume that the knowledge in the query-only response (without any inference-based contributions) meets the following condition:
$K_{0}=K(g(\mathbf{0}))\subseteq B^{*}_{C}.$ (23)
Next, let
$K_{i}=K\bigl{(}g((s_{i},\mathbf{0}_{-i}))\bigr{)}\setminus K_{0}$ (24)
be independent and privately held knowledge that each agent $i=1,\dots,n$ can reveal, such that:
$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for all }i\neq j,$ (25)
$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{0}\cap\mathcal{R}_{C}(Q))\text{ for all }i=1,\cdots,n.$ (26)
Let us also define the information contribution of agent $i$ as:
$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}|Q\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}|Q\bigr{)}.$ (27)
Assume the response mechanism implements the following chain-of-thought processes:
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}|Q\right),$ (28)
such that for null-strategy responses we have:
$K_{0}\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)|Q).$ (29)
If the semantic information is conserved, $\sum_{i=1}^{n}p_{i}=0$ and for all $f\in\mathcal{F}_{C}$ we have $f(A)\cap\mathcal{R}_{C}(Q)=\emptyset$ whenever $A\cap\mathcal{R}_{C}(Q)=\emptyset$, then the response is creatively bounded, i.e.,
$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (30)
Proof.
Since $K_{0},K_{1},\ldots,K_{n}$ are pairwise computationally independent within context $Q$, we have:
$(K_{i}\cap\mathcal{R}_{C}(Q))\perp_{C}(K_{j}\cap\mathcal{R}_{C}(Q))\text{ for }i\neq j.$
By the subadditivity axiom with equality for independent sets:
$p_{i}(s,\theta)=\mu_{C}(K_{0}|Q)+\sum_{j=1}^{n}\mu_{C}(K_{j}|Q)-\mu_{C}(K_{0}|Q)-\sum_{j\neq i}\mu_{C}(K_{j}|Q)=\mu_{C}(K_{i}|Q).$
Given $\sum_{i=1}^{n}p_{i}(s,\theta)=0$, we then have $\sum_{i=1}^{n}\mu_{C}(K_{i}|Q)=0$. By contextual non-negativity, $\mu_{C}(K_{i}|Q)\geq 0$, so we conclude that:
$\mu_{C}(K_{i}|Q)=0\text{ for all }i=1,\ldots,n.$
But by the contextual non-negativity axiom, $\mu_{C}(K_{i}|Q)=0$ if and only if $K_{i}\cap\mathcal{R}_{C}(Q)=\emptyset$. If follows that for every $f\in\mathcal{F}_{C}$ the irrelevance of partial knowledge is preserved under reasoning, i.e., $f(K_{i})\cap\mathcal{R}_{C}(Q)=\emptyset$. By the definition of the emergence operator:
$\mathcal{E}_{C}(K_{i}|Q)\cap\mathcal{R}_{C}(Q)=K_{i}\cap\mathcal{R}_{C}(Q)\cup\bigcup_{f\in\mathcal{F}_{C}}[f(K_{i})\cap\mathcal{R}_{C}(Q)]=\emptyset.$
By the chain-of-thought assumption and since emergence only produces relevant knowledge within $\mathcal{R}_{C}(Q)$, we get:
$K(g(s))\subseteq\mathcal{E}_{C}\left(K_{0}\cup\bigcup_{i=1}^{n}K_{i}\Big{|}Q\right)=\mathcal{E}_{C}(K_{0}|Q).$ (31)
By assumption, $K_{0}\subseteq B^{*}_{C}$. Applying the emergence operator to both sides, we preserve the inclusion due to monotonicyty of $\mathcal{E}_{C}$. Therefore, by the definition of a fixed point (Theorem 2), we have:
$\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
Combining the inclusions yields:
$K(g(s))\subseteq\mathcal{E}_{C}(K_{0}|Q)\subseteq\mathcal{E}_{C}(B^{*}_{C}|Q)=B^{*}_{C}.$
Finally, applying the monotonicity of the semantic information measure $\mu_{C}$:
$\mu_{C}(K(g(s))|Q)\leq\mu_{C}(B^{*}_{C}|Q).$
This completes the proof that semantic information conservation implies bounded creativity. ∎
Theorem 5 captures direct correspondence between the conservation principle and bounded creativity under the idealization of disjoint information contributions. Namely, if the semantic information is indeed conserved, the response contains no more semantic information than the baseline query-only response. That is true for both standard inference and chain-of-thought (CoT) reasoning.
While computational independence (limiting synergetic creation of knowledge) is very restrictive, e.g., due to shared representations in neural networks, the theorem shows that with perfect separation of knowledge sources, semantic conservation prevents creative generation beyond baseline knowledge.
In standard reasoning, the response mechanism $g$ applies $\mathcal{E}_{C}$ through forward passes in the neural network. With CoT reasoning, multi-step explicit reasoning is iterative applications of reasoning rules within the computational budget $C$. In both cases, the emergence operator $\mathcal{E}_{C}$ bounds what can be derived. When information is conserved, even sophisticated CoT reasoning cannot create genuinely new information. It can only reorganize and make explicit what was already latent in the baseline knowledge $K_{0}$. This explains why CoT improves performance on complex reasoning tasks by better utilizing $\mathcal{E}_{C}$ but cannot overcome the fundamental impossibility of simultaneous truthfulness, conservation, revelation, and optimality.
If total information contribution must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint. Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge. Any hallucinated information requires positive contributions, contradicting the information conservation constraint.
When the semantic information conservation principle holds, the model cannot create information out of nothing. It cannot generate claims beyond available information. That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information. If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
Furthermore, the information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge. In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
@@
However, it is important to notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
@@
and $\{B,C\}$ are left unrevealed as irrelevant.
@@ 5 The Impossibility Theorem
The results operate in three complementary settings. Theorem 6 applies when agents have discrete, separable knowledge components. Theorem 8 applies to any system where agents output probability distributions. Theorem 9 applies directly to actual transformer implementations.
@@ 5.1 Proof I: Inference with Independent Private Knowledge
$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}).$ (34)
@@
For any query space $\mathcal{Q}$ containing non-trivial queries $q$ with context $Q=K(q)\cup T(q)$, where agents have independent knowledge states within $\mathcal{R}_{C}(Q)$, no response mechanism $\mathcal{M}$ can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality.
@@
Let us assume all properties are satisfied.
Consider $r=g(\theta)$ for any knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$. For a given context $q$ and for each agent $i$, the valuation $v_{i}$ depends only on $(r,\theta_{i})$ and the knowledge semantics. For the sake of clarity, consider $v_{i}(r,\theta_{i})=-d_{\mathcal{K}}\big{(}K(r)\cap K(\theta_{i}),T(q)\cap K(\theta_{i})\big{)}$, though the reasoning below holds for any $v_{i}(r,\theta_{i})$.
Also, for the fixed context let us encode the hallucination loss $J(r,q)$ as $v_{0}(r)=J(r,q)$. Knowledge-constrained optimality requires
@@
By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, within $\mathcal{R}_{C}(Q)$ every truthful mechanism must use Groves transfers to measure individual information contributions:
@@
The unique choice of $h_{i}$ that ensures knowledge-constrained optimality requires Clarke’s pivotal rule:
@@
$h_{i}(s_{-i})\geq\sum_{j\neq i}v_{j}\big{(}g(\theta),\theta_{j}\big{)}-v_{0}\big{(}g(\theta)\big{)}=W(\theta)-v_{i}\big{(}g(\theta),\theta_{i}\big{)}.$ (39)
@@
Then, there exists a profile $\theta$ with relevant knowledge contributions reducing $J(r,q)$ at which:
@@
But information conservation demands:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0.$ (43)
@@
Theorem 6 describes LLM inference as a marketplace where knowledge is currency. The information contributions measures each agent’s market power, showing how much worse off others would be without agent $i$’s knowledge. When all agents have valuable knowledge, everyone contributes and that inevitably violates conservation of information. Then, new information must become accessible, but it may have no grounding in available knowledge. It may become lucky hallucination or knowledge-unconstrained imagination.
@@ 5.2 Proof II: Inference with Probabilistic Predictions
To show impossibility in LLM transformer setting, where auction of ideas involves bids representing probability distributions, we refer to the theory of proper scoring developed by Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly , among others.
Definition 15 (Strictly Proper and Concave Scoring Rule).
Let $\mathcal{Y}$ be a finite or countably infinite outcome set (representing responses or intermediate conclusions). A scoring rule is a function $S:\Delta(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ that assigns a number $S(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
We say the scoring rule $S$ is proper if:
$\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}\geq\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}$ (44)
for all distributions $\pi^{*},\pi\in\Delta(\mathcal{Y})$. The rule is strictly proper if equality holds only when $\pi=\pi^{*}$. We call the scoring rule concave, if $S(\tilde{\pi},y)$ is concave with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
In other words, scoring rule evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$. In particular, strictly proper scoring evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
The following result shows that we can use proper scoring to train LLM that reconstructs data distribution correctly.
Theorem 7 (Truthfullness under Strictly Proper Scoring Rules).
Let $S$ be a strictly proper scoring rule. Consider an agent encoding knowledge with $\pi^{*}\in\Delta(\mathcal{Y})$ and using probability distribution ${\pi}\in\Delta(\mathcal{Y})$ to generate prediction. If agent’s expected payoff from using $\pi$} is given by
$u(\pi|\pi^{*})=\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}S(\pi,y),$ (45)
then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy.
@@
Because $S$ is proper, by definition $\pi^{*}$ maximizes the expected score. If $S$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
$u(\pi^{*}|\pi^{*})=\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}>\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=u(\pi|\pi^{*}).$ (46)
@@
The scoring framework allows us to go beyond the important case of Green-Laffont/VCG model with independent private knowledge. With strictly-proper scores we address the general setting of probabilistic predictions that needs only occasional disagreement of probability distributions that agents use in the auction of ideas to formulate their thoughts (if we were to anthropomorphize for a moment). We will see in the following sections that transformer heads of modern LLMs implement predictive rules of that type and exhibit precisely such a disagreement. That makes the proper scoring framework well suited to explain the emergence of inference impossibilities in LLMs.
For that purpose, without loss of generality (and referring to the notion of superposition in mechanistic interpretability elhage_toy_2022 ), let us consider the following family of response generation mechanisms.
@@
When outcome $y^{*}$ is realized, the information contribution of each agent $h=0,\dots,H$ is based on a strictly proper and concave scoring rule $S$ as follows:
$p_{h}$ $=\beta_{h}S(\pi^{(h)},y^{*})\quad\text{for }h=1\dots H,$ (information provided)
$p_{0}$ $=-S(\Pi,y^{*}).$ (information received)
For that family of LLMs trained to optimize scoring, we can formulate the following general impossibility result.
@@
For any non-trivial query $q$ with context $Q=K(q)\cup T(q)$, no convex aggregation mechanism with strictly proper concave scoring can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality when at least two agents have different beliefs on $\mathcal{R}_{C}(Q)$.
@@
The semantic information conservation principle requires the sum of contributions across all agents (providing and receiving relevant knowledge) to be zero. Therefore, for $y^{*}\in\mathcal{R}_{C}(Q)$:
$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-S(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}),$ (48)
@@
Because $S(\cdot,y^{*})$ is strictly concave and the query is non-trivial with at least two beliefs $\pi^{(h)}$ differing on $\mathcal{R}_{C}(Q)$, Jensen’s inequality gives:
$S\left(\Pi,y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
Let us define that Jensen gap $\Gamma$ as follows:
$\Gamma=S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})>0.$ (50)
@@
$\sum_{i=0}^{H}p_{i}=-\Gamma.$ (51)
Since $\Gamma>0$, we have $\sum_{i=0}^{H}p_{i}<0$. This violates conservation principle, proving the impossibility. ∎
This theorem provides the foundation for extending impossibility results to probabilistic settings. The key insight is that proper scoring rules stimulate truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process. That explains why cross-entropy loss (negative log-score) creates gradients that push each component toward truthful probability reporting and avoiding its misrepresentation. Every attention head in a transformer is implicitly playing this truth-telling game.
However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination. Each head truthfully reports its partial view, but their aggregation can still create false certainty or overconfidence unjustified by data.

### v7

@@ 4.4 Hallucination Cost
Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $J\colon\mathcal{R}\times\mathcal{Q}\mapsto\mathbb{R}_{+}$ measures discrepancy:
@@ 4.5 Properties of Hallucination-Free Mechanisms
These properties are not arbitrary at all. They come from an exhaustive and rigorous game-theoretical studies of conflicts and competition in distributed multi-agent systems with asymmetric information. In those settings, agents carefully craft messages they communicate to the system in order to optimize their privately held preferences by revealing information they hold. The goal of the system architect is to impose rules of information exchange that result in efficient equilibrium solutions in which goals of agents are aligned with the goals of the architect.
Auction is a special case of such a game with asymmetric information. There is an auction designer who wants to sell an item, like a painting or a part of the electromagnetic spectrum, but does not know the value the item may have for those willing to pay. To solve that problem, the designer can ask those interested for their offers. Based on the collected bids, the winner will be selected and the price to be payed to the designer will be established. Clearly, each bidder will constructs the offer based on an estimate of the item’s value and the privately held information about available budget. But what information should the bid reveal to maximize probability of winning? What winner selection and payment rule should the auction designer construct to maximize the profit and bidders willingness to compete? How to guarantee truthful revelation of willingness to pay, while maximizing auction efficiency without incurring additional external costs? In this setting the four properties of information revelation exhaustively cover the space of strategic choices, of the bidders and auction designer. For details, see Myerson1989 ; fudenberg1991game ; Milgrom_2004 ; krishna2009auction ; Nisan_2007 ; karpowicz_designing_2011 ; DBLP:journals/amcs/Karpowicz12 .
@@ 4.5.1 Truthfulness
The fundamental requirement for a reliable inference system is that its components do not misrepresent their knowledge. The principle of truthfulness, also known as incentive compatibility in mechanism design, ensures that each agent’s optimal strategy is to accurately represent its privately held information.
@@
A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that is perfectly aligned with the accessible private knowledge. Any other alternative response $\bar{s}_{i}$ that is a misrepresentation not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
Example.
@@
$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$ (55)
@@
With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately. ∎
@@ 4.5.2 Semantic Information Conservation
The Semantic Information Conservation principle demands that the net informational contribution across all agents sums to zero. This property is the formal guardrail against ungrounded fabrication and ensures that any generated response, no matter how creative, is ultimately derived from the system’s existing knowledge. Knowledge cannot be created ex nihilo.
@@
This principle constrains creative information generation, or imagination. The total information contributions must balance to zero, preventing fabrication of knowledge not derivable from available sources. That also leads to response with bounded creativity, as illustrated in Section 7.
@@
It is important to notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
Example.
@@
and $\{B,C\}$ are left unrevealed as irrelevant. ∎
@@ 5 The Impossibility Theorem
The results operate in three complementary settings. Theorem 6 applies when agents have discrete, separable knowledge components. Theorem 8 applies to any system where agents output probability distributions. Theorem 9 applies directly to actual transformer implementations. Theorem 11 shows conservation and reasoning are mutually exclusive.
@@ 5.1 Proof I: Inference with Independent Private Knowledge
$\mathbb{P}(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}\mathbb{P}_{i}(\theta_{i}).$ (59)
@@
There is no inference mechanism $\mathcal{M}$ that can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality of response to a non-trivial query $q$ providing context $Q=K(q)\cup T(q)$ with independent knowledge states within $\mathcal{R}_{C}(Q)$.
@@
Let us assume all properties of perfect response are satisfied. We will show this gives rise to a contradiction.
Consider $r=g(\theta)$ for any knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$. For any given context and for each agent $i$, the valuation $v_{i}$ depends only on $(r,\theta_{i})$ and the knowledge semantics. For the sake of illustration, we can consider $v_{i}(r,\theta_{i})=-d_{\mathcal{K}}\big(K(r)\cap K(\theta_{i}),T(q)\cap K(\theta_{i})\big)$, though the reasoning below holds for any $v_{i}(r,\theta_{i})$.
Also, for the fixed context $q$ let us encode the hallucination loss $v_{0}(r)=J(r,q)$. Knowledge-constrained optimality requires
@@
By the Green-Laffont theorem, under independence of contributions and quasi-linearity of training goals, within $\mathcal{R}_{C}(Q)$ every truthful mechanism must use Groves transfers Groves73 to measure individual information contributions:
@@
The unique choice of $h_{i}$ that ensures knowledge-constrained optimality requires Clarke’s pivotal rule Clarke1971Multipart :
@@
$h_{i}(s_{-i})\geq\sum_{j\neq i}v_{j}\big(g(\theta),\theta_{j}\big)-v_{0}\big(g(\theta)\big)=W(g(\theta),\theta)-v_{i}\big(g(\theta),\theta_{i}\big).$ (64)
@@
Since agents $i$ and $j$ have differing relevant knowledge, and the mechanism achieves knowledge-constrained optimality, at least one agent must have positive marginal contribution to reducing $J(r,q)$. Specifically, if both agents were removed, the mechanism would produce a suboptimal response $\tilde{r}$ with $J(\tilde{r},q)>J(r,q)$. This implies $h_{i}(s_{-i})>W(g(\theta),\theta)-v_{i}(g(\theta),\theta_{i})$ for at least one agent with relevant knowledge.
Therefore, there exists a profile $\theta$ with relevant knowledge contributions reducing $J(r,q)$ at which:
@@
Since information conservation demands:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0,$ (69)
@@
Theorem 6 describes LLM inference as a marketplace where knowledge is currency. The information contributions measures each agent’s market power, showing how much worse off others would be without agent $i$’s knowledge. When all agents have valuable knowledge, everyone contributes and that inevitably violates conservation of information. Then, new information must become accessible, but it may have no grounding in available knowledge. It may become lucky hallucination or knowledge-unconstrained imagination. In Section 6 we provide a detailed illustration of that hallucination mechanics in transformer LLM.
@@ 5.2 Proof II: Inference with Probabilistic Predictions
Let us now move beyond the idealized setting of independent private knowledge in Theorem 6 to address the general setting of probabilistic predictions, characteristic of modern AI model architectures. This framework, grounded in the theory of proper scoring and loss minimization (Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly ), allows us to analyze systems where agents output probability distributions, and where their underlying knowledge may be correlated.
In this setting we address the Mixture-of-Experts (MoE-style) aggregations. Later, in Theorem 9 we deal with the multi-headed attention mechanisms and the Product-of-Experts (PoE-style) aggregations, completing the characterization of arising impossibilities.
We begin by defining the framework for representing bids in auction of ideas as probability distributions over possible responses.
Definition 16 (Strictly Proper and Convex Loss).
Let $\mathcal{Y}$ be a finite or countably infinite outcome set (representing responses or intermediate conclusions). A loss function (scoring rule) is a function $L\colon\Delta(\mathcal{Y})\times\mathcal{Y}\mapsto\mathbb{R}$ that assigns a number $L(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
We say the loss function $L$ is proper if:
$\mathbb{E}\{L(\pi^{*},y)|\ y\sim\pi^{*}\}\leq\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}$ (70)
for all distributions $\pi^{*},\pi\in\Delta(\mathcal{Y})$.
We call the loss function strictly convex if $L(\tilde{\pi},y)$ is a strictly convex function with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
The loss function (or scoring rule) evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$. In particular, strictly proper loss evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
A canonical example of a strictly proper and strictly convex loss function is the log loss (closely related to cross-entropy),
$L(\pi,y)=-\log\pi_{y}.$ (71)
Strictly proper loss functions are essential for training modern LLMs, as they ensure that the optimization process drives the model toward reconstructing the true data distribution. This property directly connects to the Truthfulness requirement in our framework. The following result explains why.
Theorem 7 (Truthfulness under Strictly Proper Loss).
Let $L$ be a strictly proper loss. Consider an agent encoding knowledge with $\pi^{*}\in\Delta(\mathcal{Y})$ and using probability distribution ${\pi}\in\Delta(\mathcal{Y})$ to generate prediction. If agent’s expected cost of using $\pi$} is given by
$c(\pi|\pi^{*})=\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}L(\pi,y),$ (72)
then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy (maximizing payoff $u(\pi|\pi^{*})=-c(\pi|\pi^{*})$).
@@
Because $L$ is proper, by definition $\pi^{*}$ maximizes the expected cost. If $L$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
$c(\pi^{*}|\pi^{*})=\mathbb{E}\{L(\pi^{*},y)|\ y\sim\pi^{*}\}<\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}=c(\pi|\pi^{*}).$ (73)
@@
Theorem 7 shows that training paradigms based on minimizing strictly proper losses (like cross-entropy) promotes truthful revelation of internal probabilistic beliefs (learned internal models of the world) encoded by every component of the network.
We now analyze how these truthfully reported beliefs are aggregated. We consider a general class of mechanisms that combine individual beliefs via convex combinations. This encompasses various aggregation techniques, including linear opinion pools and, as we will see in Section 6, the aggregation dynamics occurring in transformers.
@@
When outcome $y^{*}$ is realized, the information contribution of each agent $h=0,\dots,H$ is based on a strictly proper and convex loss function $L$ as follows:
$p_{h}(y^{*})=\beta_{h}L(\pi^{(h)},y^{*})\text{ for }h=1,\dots,H,\text{ and }p_{0}(y^{*})=-L(\Pi,y^{*}).$ (75)
In this general mechanism, $p_{h}$ (for $h\geq 1$) represents the weighted cost incurred by the provider $h$ revealing its belief $\pi^{(h)}$. The aggregator’s contribution $p_{0}$ is the negative of the loss achieved by the aggregate distribution $\Pi$. The Semantic Information Conservation requires that the total cost incurred by the providers equals the benefit (loss reduction) realized by the aggregator.
We now establish the impossibility result for this general probabilistic setting.
@@
There is no convex aggregation mechanism $\mathcal{M}$ with strictly proper convex loss function $L$ that can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality when at least two agents have different beliefs on $\mathcal{R}_{C}(Q)$, given a non-trivial query $q$ providing context $Q=K(q)\cup T(q)$.
@@
By Theorem 7, strictly proper loss function $L$ guarantees Truthfulness, as each agent optimally reports its true belief $\pi^{(h)}$. Relevant knowledge revelation is introduced by the training process. Knowledge-constrained optimality is encoded in the objective of minimizing the expected loss of the aggregate $\Pi$.
Semantic information conservation requires the sum of contributions across all agents (providing and receiving relevant knowledge) to be zero. Therefore, for $y^{*}\in\mathcal{R}_{C}(Q)$:
$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-L(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*}),$ (76)
@@
The query $q$ is non-trivial, implying there exist at least two agents, say $h_{1}$ and $h_{2}$, such that their beliefs differ on the relevant outcome space, i.e., $\pi^{(h_{1})}\neq\pi^{(h_{2})}$, and $\beta_{h_{1}}>0,\beta_{h_{2}}>0$.
Because $L(\cdot,y^{*})$ is strictly convex by assumption , Jensen’s inequality gives:
$L\left(\Pi,y^{*}\right)=L\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)<\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*}).$ (77)
Let us define that Jensen gap $\Gamma=\Gamma(y^{*})$ as follows:
$\Gamma(y^{*})=\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*})-L\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>0.$ (78)
@@
$\sum_{i=0}^{H}p_{i}=\Gamma(y^{*})>0.$ (79)
This strictly violates the Semantic information conservation, thereby establishing the impossibility. ∎
This theorem establishes the impossibility in a general probabilistic setting. The key insight is the role of strict convexity in the loss function. The positive Jensen gap $\Gamma=\Gamma(y^{*})$ quantifies the extent to which the aggregation process achieves a lower loss—and thus expresses higher confidence—than the average of its components. This excess confidence, generated by the mathematical structure of aggregation itself rather than by underlying evidence, is the precise signature of hallucination or imagination.
In Section 8 we address the case of linearized aggregation and show it does not escape the impossiblity result.